OpenStudy (anonymous):

Solve the equation for y: 2x + 3(y + 9) = 17.

OpenStudy (nali):

first multiply 3 times y+9 using distitributive property

OpenStudy (nali):

after that move the 2x to the other side by subtracting 2x form both sides

OpenStudy (nali):

at the end just devide by 3 to get the y by it self and ur done

OpenStudy (anonymous):

so what should be the final answer

OpenStudy (nali):

well i basically give u the final answer when i told u the exact steps plz do it first and show me that u understand what i stated above

OpenStudy (anonymous):

i did it i just wanna make sure that i did it right i got y = 5-2x

OpenStudy (nali):

ok let me check

OpenStudy (nali):

no that is not right

OpenStudy (anonymous):

well than what is

OpenStudy (nali):

y=-2x/3 -10/3

OpenStudy (nali):

by the way there is supposed to be a step between my first and my second comment it is this then moove the 27 to the other side by subtracting 27 form both sides

OpenStudy (nali):

@crhoades96 did u get the answer and how i got to it ?

OpenStudy (anonymous):

2x+3y+18=17 -18 both side 2x+3y=-1 3y=-1+2x

OpenStudy (nali):

@lovemie ur first step is wrong 3 times 9 is not 18 it is 27 that is the same mistake i did before

OpenStudy (anonymous):

omg yeah it's 27

OpenStudy (amistre64):

there really is no need to distribute the 3 thru; its harmless, but unnecessary